OpenStudy (anonymous):

using Euclidean algorithm find GCD(1071,462)

OpenStudy (zzr0ck3r):

1071 = 462*2+147 462 = 147*3+21 147 = 21*7+0 so the answer is 21

OpenStudy (kc_kennylau):

\[\begin{array}{rcl} 1071&=&2\times\color{blue}{462}&+&\color{green}{147}\\ \color{blue}{462}&=&3\times\color{green}{147}&+&\color{red}{21}\\ \color{green}{147}&=&7\times\color{red}{21}&+&0 \end{array}\]Therefore GCD(1071,462) is 21

OpenStudy (zzr0ck3r):

so you do that process until you get 0 as a remainder then the answer is the last remainder you were given

OpenStudy (zzr0ck3r):

very pretty

OpenStudy (zzr0ck3r):

and now you can find the LCM very easy as well:)

OpenStudy (kc_kennylau):

@sherahhhh what do you not understand? :)